# Cosmology using numerical relativity

### v1

@@ 2.7 Units
Cosmologists use units in which $\hbar=c=1$, and quantities are described in terms of the reduced Planck mass $M_{\mathrm{Pl}}=\sqrt{\hbar c/8\pi G}$. In this system of units masses $\sim M_{\mathrm{Pl}}$ and lengths $\sim 1/M_{\mathrm{Pl}}$ , unlike in NR where $G=c=1$ and both lengths and masses have units of $[M]$ where $M$ is some physical mass scale in the problem (e.g. $M=M_{\odot}$). So, for example, energy density has units of $[M_{\mathrm{Pl}}]^{4}$ in cosmology and $[M]^{2}$ in NR. To convert a quantity X with units of $[M]^{n}$ in NR units to its cosmological value Y in units of $[M_{\mathrm{Pl}}]^{m}$, one uses
@@ 3.2 Inflation and the initial condition problem
Cosmic inflation [132, 133, 134, 135, 136] is a theory of the early universe in which a period of nearly exponential expansion occurs, providing an explanation of why the universe on large scales is spatially flat and homogeneous, despite the expectation that the cosmological singularity will be chaotic and, due to its local nature, uncorrelated at separated points. Inflation also provides a natural mechanism to generate an approximately scale-invariant spectrum of primordial perturbations that seed structure formation, as observed in the CMB [137]. The simplest mechanism for inflation is a single scalar field, the inflaton, that slowly rolls down a roughly flat region of its potential. To be consistent with this picture, sufficiently generic (presumably inhomogeneous) initial data should still permit the scalar field potential to dominate and at least some of the universe to undergo a period of inflation. Single-field models of inflation are classified as high-scale or low-scale models, depending on the energy scale associated with inflation, or as large-field or small-field models depending on the field range over which the minimum required 60 e-folds of inflation occurs (some care is needed in interpreting the different terms [138]). In cases where the inhomogeneities are strongly backreacting at the Hubble scale (which is predominantly the case for large-field models), NR simulations may be needed to determine whether this happens.
@@
For high-scale inflation, a number of simulations have supported the claim that inflation persists with generic initial conditions. The first 3+1D simulations with periodic boundary conditions by Laguna, Kurki-Suonio & Matzner [143, 144] showed that inflation could occur even if kinetic and gradient energies were initially comparable to the potential energy density. However, these simulations were limited to initial conditions that did not lead to BH formation. Simulations leading to BH formation were carried out by Shibata et al. [54] for axisymmetric spacetimes with a cosmological constant and an initially large distribution of gravitational waves. While their boundary conditions (asymptotically de Sitter) and the presence of a cosmological constant meant that the end state is always de Sitter [145], they showed that there can be an initial gravitational wave radiation dominated period (i.e. the GW do not always collapse to form a Schwarzschild-de Sitter spacetime.) It would therefore be interesting to revisit the work with a more general inflationary model and/or boundary setup. More general 3+1D initial conditions, including those that lead to BH formation, were considered first in East et al. [90] and studied later by others [91, 92, 146, 93, 80, 72, 75]. These showed that high-scale inflation can occur in the presence of large inhomogeneities even where BHs form, see Fig. 9. Some recent work by Garfinkle, Ijjas & Steinhardt has opposed this conclusion on the basis that the initial conditions and boundary choices made may bias the results towards successful inflation in some non-generic regions of low initial curvature [49, 147], see Joana [148] for a partial response.
@@
The effect of large non-uniform initial inflaton velocity in addition to an inhomogeneous initial field profile was first studied in Corman et al. [72] for the large-field case in both single and two-field inflationary models, and it was found that these models are generally robust, a conclusion that was later supported by the work of Elley et al. [75] using different methods for generating the initial conditions. In Corman et al. [72], the XCTS method with a spatially constant mean curvature $K$ and a rescaled scalar field ansatz were used to generate initial data, as well as an ansatz for the momentum constraint suggested by Garfinkle & Mead [67], whereas the latter paper used the CTTK method of Aurrekoetxea et al. [48] that creates a spatially varying mean curvature $K$ and greater gradients in the extrinsic curvature $A_{ij}$. Despite the different initial data choices, and the use of GHC versus BSSN formulations, both works showed that the effect of velocity on low-scale models resulted in a larger spread of e-folds in different regions [75], but did not significantly improve or harm the success of such models, in comparison to the zero velocity case – see Fig. 10. This work addressed the concerns of Ijjas [13] regarding the effect of kinetic inhomogeneities (but does not respond to the more recent criticisms in [49, 147]).
@@ 3.3 Ekpyrotic and bouncing scenarios
This string-inspired model has developed into a more general paradigm in which the universe undergoes cyclic periods of expansion and collapse, separated by a non-singular bounce [155, 156]. In such cases the slow contraction phase leading to smoothing of inhomogeneities, rather than an exponential expansion as in inflation (Ijjas & Steinhardt provide a pedagogical review in [157]).
@@
to the usual slow contraction field $\phi$. This ghost field violates the Null Energy Condition, but since its energy density grows negative faster than the positive energy of $\phi$ as the scale factor $a(t)\rightarrow 0$, it will dominate at small $a(t)$ and turns the contraction phase into an expansion phase, executing the bounce [170]. Xue et al. [169] studied the evolution of perturbations through the bounce using the harmonic gauge (see Sec. 2.5), and found that if the initial perturbations are large, the induced non-linearities can cause inhomogeneous regions to collapse into a singularity while relatively homogeneous regions will undergo the bounce successfully. In a slightly different direction Corman et al. [53] studied the behaviour of BHs through such a two-field bounce model, and found that for sufficiently large BHs (with horizons $\sim 3-4$ times the order of the minimum cosmological horizon) the two horizons can merge and disappear for a short time during the contracting phase. Nevertheless, this state of affairs is temporary and as the bounce enters the expanding phase, the horizons separate and the BH end state is similar to its initial pre-contraction phase value.
Figure 12: (Left:) Figure from Ijjas et al. [168] showing slow contraction in the multi-field case needed to generate scale invariant perturbations. Time evolution of the normalized energy densities in the fields (green) the curvature (red) and the shear (blue) are shown. The model drives the universe towards FRW, but a non-zero gradient in the auxilliary field destabilizes the FRW state on longer timescales. (Right:) Figure from Xue et al. [169] showing the local expansion as a function of space, evolving over time. Whilst the whole of the space bounces (going from negative to positive expansion), the space in the middle bounces at a much later time compared to the other regions, creating some inhomogeneity at late times.
@@ 3.4 Signatures of inflationary bubble collisions
Figure 13: Schematic figure depicting the procedure from a scalar inflationary potential to a prediction of the observable comoving curvature perturbation $\mathcal{R}$. NR simulations are performed to construct the full collision spacetime (upper right). The perturbed FRW metric inside the observation bubble is reconstructed by evolving a set of geodesics through the simulation (lower right panel). A gauge transformation is used to extract the comoving curvature perturbation $\mathcal{R}$ late in the inflationary epoch. Evolving the comoving curvature perturbation, the observable universe is split into a region that is, and a region that is not, affected by the collision. The lower left panel depicts the surface of last scattering inside the observation bubble in a reference frame where the observer is at the origin of coordinates. Figure from Wainwright et al. [194].
@@ 4.1 Reheating and oscillon formation
Another example of a preheating process is gauge preheating, when the energy of the inflaton is transferred into gauge fields. This can be sourced via a coupling $\alpha_{g}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$ in the Lagrangian, where $\alpha_{g}$ parameterizes the strength of the coupling and the range of momenta for which the process features a tachyonic instability. This instability allows the full vacuum energy to be transferred within a single oscillation of the inflaton field, and the strongest effects occur when the inflaton rolls the fastest ($\dot{\phi}$ is the largest) – typically at the end of inflation. In Adshead et al. [102], the authors showed that linearized gravity is quickly violated during this process. However, despite reaching large density contrasts $\delta\rho/\rho\approx 30$, PBHs did not form, indicating that the internal pressure of the gauge fields was playing an important role. Using a similar approach to Aurrekoetxea et al. [73] but based on the over-mass $\delta M$ enclosed in a region of radius $R$, they computed the compactness of the overdensities formed and found a maximum value of $\mathcal{C}\approx 0.1$. The authors compared the extracted GW spectrum from FLRW and BSSN lattice simulations finding remarkable agreement and suggesting that non-linear gravity at most had a $\mathcal{O}(1)$ effect, see Fig. 17.
@@ 4.2 Phase transitions and topological defects
As the universe expands and cools down, it can undergo phase transitions that lead to the formation of exotic objects. Among the most studied relics are topological defects [209, 210, 211] – structures formed when the universe experienced symmetry breaking. The specific details of these structures depend on the topology of the vacuum. For example, if the vacuum breaks a parity symmetry, two-dimensional defects known as domain walls can form. We previously encountered such relics in Sec. 3.4, where we discussed the observational signatures in the CMB from inflationary bubble collisions. NR has also been used to study the evolution of individual bubbles during vacuum decay, such as the 1+1D work by Blanco-Pillado et al. [212], and Giombi et al. [213]. Recently, Lin et al. [214, 215] have performed 3+1D NR simulations to investigate how the evolution of inhomogeneities in the early universe can populate the landscape of multiple vacua. The dynamics of the formed bubbles may serve as seeds for supermassive black holes, as studied by Huang et al. [216], and enhance the formation of black hole binaries, as suggested by Yuwen et al. [217].
@@
The dynamics of topological defects sources source gravitational waves. If these are weak, we might observe the incoherent sum of all the events that are too weak to be directly detected – the stochastic GW background. However, local dynamics might also source strong gravitational-wave signatures that can be directly detected. In both cases, an NR treatment may be necessary: either to validate the predictions of a weak-field treatment of the source, or to predict the detailed signatures to be expected. However, much of the current research focuses on the linearised treatment of these gravitational-wave signatures, and to the best of our knowledge, there is no simulation that fully incorporates NR in a cosmological context – hence the brevity of this section. However, asymptotically flat NR simulations have been performed to understand the GR dynamics of these objects and construct gravitational-wave templates [220, 221, 222, 223, 224].
@@ 4.3 Primordial black holes
Using spheroidal super-horizon perturbations, Yoo et al. [266] studied the effect of ellipticity on the threshold of PBH formation, finding that it is negligible in standard situations during the radiation-dominated era. Yoo later extended the work [267] by adding a tidal torque to the non-spherical profile of the initial curvature perturbation as a mechanism to generate angular momentum and form a spinning remnant, but the final spin remained consistent with zero, in agreement with expectations for the radiation-dominated era. However, in the case of a softer equation of state, small deviations from spherical symmetry may evolve differently due to the absence of pressure gradient forces. De Jong et al. [268, 74] used a simple model of two scalar fields to study the mass and spin of PBHs from the collapse of perturbations with angular momentum during an early matter-dominated era $P\approx 0$. They found that although the initial mass and spin at formation can be of order $\mathcal{O}(10\%)$, the accretion of background matter post-collapse dilutes the dimensionless spin of the PBH. Hence, unless the matter era is short, they concluded that the final spin of the PBHs formed will be negligible.

### v2

@@ 2.7 Units
Cosmologists often use units in which $\hbar=c=1$ for studies of the early universe, where quantities are described in terms of the reduced Planck mass $M_{\mathrm{Pl}}=\sqrt{\hbar c/8\pi G}$. In this system of units masses $\sim M_{\mathrm{Pl}}$ and lengths $\sim 1/M_{\mathrm{Pl}}$ , unlike the geometric units used in most simulations in NR where $G=c=1$ and both lengths and masses have units of $[M]$ where $M$ is some physical mass scale in the problem (e.g. $M=M_{\odot}$). So, for example, energy density has units of $[M_{\mathrm{Pl}}]^{4}$ in Planck units and $[M]^{2}$ in geometric units. Another important case to highlight is that the scalar field is dimensionless in geometric units, as opposed to having a mass dimension in Planck units. A value of 1.0 for the scalar field is therefore always very large and corresponds to a Planckian energy scale, regardless of what physical scale [M] is taken to represent. To convert a quantity X with units of $[M]^{n}$ in geometric units to its value Y in units of $[M_{\mathrm{Pl}}]^{m}$, one uses
@@ 3.2 Inflation and the initial condition problem
Cosmic inflation [179, 180, 181, 182, 183] is a theory of the early universe in which a period of nearly exponential expansion occurs, providing an explanation of why the universe on large scales is spatially flat and homogeneous, despite the expectation that the cosmological singularity will be chaotic and, due to its local nature, uncorrelated at separated points. Inflation also provides a natural mechanism to generate an approximately scale-invariant spectrum of primordial perturbations that seed structure formation, as observed in the CMB [184]. The simplest mechanism for inflation is a single scalar field, the inflaton, that slowly rolls down a roughly flat region of its potential. To be consistent with this picture, sufficiently generic (presumably inhomogeneous) initial data should still permit the scalar field potential to dominate and at least some of the universe to undergo a period of inflation. Single-field models of inflation are classified as high-scale or low-scale models, depending on the energy scale associated with inflation, or as large-field or small-field models depending on the field range over which the minimum required 60 e-folds of inflation occurs (some care is needed in interpreting the different terms [185]). In cases where the inhomogeneities are strongly backreacting at the Hubble scale (which is predominantly the case for large-field models), NR simulations may be needed to determine whether this happens [^15] Early work on using NR techniques to investigate stochastic inflation has been done by Launay, Rigopoulos & Shellard [186] and Florio & Shellard [187]..
@@
For high-scale inflation, a number of simulations have supported the claim that inflation persists with generic initial conditions. The first 3+1D simulations with periodic boundary conditions by Laguna, Kurki-Suonio & Matzner [192, 193], building on earlier 1D work with Centrella and Wilson [12], showed that inflation could occur even if kinetic and gradient energies were initially comparable to the potential energy density. However, these simulations were limited to initial conditions that did not lead to BH formation. Simulations leading to BH formation were carried out by Shibata et al. [87] for axisymmetric spacetimes with a cosmological constant and an initially large distribution of gravitational waves. While their boundary conditions (asymptotically de Sitter) and the presence of a cosmological constant meant that the end state is always de Sitter [194], they showed that there can be an initial gravitational wave radiation dominated period (i.e. the GW do not always collapse to form a Schwarzschild-de Sitter spacetime.) It would therefore be interesting to revisit the work with a more general inflationary model and/or boundary setup. Another early work in 1D planar symmetry showed that GWs cannot prevent inflation with a cosmological constant or massive field [13, 14]. More general 3+1D initial conditions, including those that lead to BH formation, were considered first in East et al. [130] and studied later by others [131, 132, 195, 133, 119, 109, 113]. These showed that high-scale inflation can occur in the presence of large inhomogeneities even where BHs form, see Fig. 9. Some recent work by Garfinkle, Ijjas & Steinhardt has opposed this conclusion on the basis that the initial conditions and boundary choices made may bias the results towards successful inflation in some non-generic regions of low initial curvature [80, 196], see Joana [197] for a partial response. The key issue that needs to be addressed is whether the observations of successful inflation in other studies arise only from those parts of the spacetime which were “non generic” - i.e. the regions for which the initial curvatures were forced to be low by the imposition of periodic boundary conditions. In the previous simulations, large curvature regions often collapsed to BHs, but for the reasons discussed in the introductory sections of this review, directly tracking the causally related parts of the spacetime is non trivial and cannot be done simply by comparing aligned coordinate regions, as was done in [80]. It should, however, be possible to consider only the regions causally connected to a region of high curvature (by tracking null geodesics from the initial slice, for example), and then evaluate whether any of this restricted region undergoes inflation. If not, and the whole region ends up in a black hole, this would potentially validate the concerns expressed in [80]. However, in at least some scenarios that were considered in the earlier works [130, 131, 132, 195, 133, 119, 109, 113] black holes did not form and the final spacetime was fully inflationary, so at least in such cases it seems likely that inflation would survive, albeit with a limit on the magnitude of initial perturbations.
@@
The effect of large non-uniform initial inflaton velocity in addition to an inhomogeneous initial field profile was first studied in Corman et al. [109] for the large-field case in both single and two-field inflationary models, and it was found that these models are generally robust, a conclusion that was later supported by the work of Elley et al. [113] using different methods for generating the initial conditions. In Corman et al. [109], several methods based on CTS with a spatially constant mean curvature $K$ were used to generate initial data – firstly choosing a constant time derivative for the scalar field, and secondly by rescaling the momentum such that the ratio of the gradient and kinetic energy after solving for the constraints is constant, which allowed them to look at scenarios where there is the same amount of energy in the gradient and kinetic terms. They also used a rescaled scalar field ansatz for the momentum constraint suggested by Garfinkle & Mead [104]. The paper by Elley et al. [113] used the CTTK method of Aurrekoetxea et al. [79] that creates a spatially varying mean curvature $K$ and greater gradients in the extrinsic curvature $A_{ij}$. Despite the different initial data choices, and the use of GHC versus BSSN formulations, both works showed that the effect of velocity on low-scale models resulted in a larger spread of e-folds in different regions [113], but did not significantly improve or harm the success of such models, in comparison to the zero velocity case – see Fig. 10. This work addressed the concerns of Ijjas [33] regarding the effect of kinetic inhomogeneities (but does not respond to the more recent criticisms in [80, 196]).
@@ 3.3 Ekpyrotic and bouncing scenarios
This string-inspired model has developed into a more general paradigm in which the universe undergoes cyclic periods of expansion and collapse, separated by a non-singular bounce [204, 205]. In such cases the slow contraction phase leading to smoothing of inhomogeneities, rather than an exponential expansion as in inflation (Ijjas & Steinhardt provide a pedagogical review in [206]). It should be noted that to subsequently transition from a state of collapse to a state of expansion - i.e. to achieve a “bounce” - one needs to violate the null convergence condition (NCC), which in GR is equivalent to violating the null energy condition (NEC). This is a stronger condition than the violation of the strong energy condition (SEC) required for an inflationary fluid, and whilst the energy conditions are known to be violated in certain cases (e.g. the Casimir effect), violating the NEC on large scales requires some new physics or exotic matter to come into play.
@@
to the usual slow contraction field $\phi$. This ghost field violates the Null Energy Condition, but since its energy density grows negative faster than the positive energy of $\phi$ as the scale factor $a(t)\rightarrow 0$, it will dominate at small $a(t)$ and turns the contraction phase into an expansion phase, executing the bounce [219]. Xue et al. [218] studied the evolution of perturbations through the bounce using the harmonic gauge (see Sec. 2.5), and found that if the initial perturbations are large, the induced non-linearities can cause inhomogeneous regions to collapse into a singularity while relatively homogeneous regions will undergo the bounce successfully. In a slightly different direction Corman et al. [85] studied the behaviour of BHs through such a two-field bounce model with black hole horizons ranging from 1/8th the size of the minimum cosmological horizon (if there was no black hole) to order 1, and found that for sufficiently large BHs, the apparent horizons can merge and disappear for a short time during the contracting phase. Nevertheless, this state of affairs is temporary and as the bounce enters the expanding phase, the horizons separate and the BH end state is similar to its initial pre-contraction phase value.
Figure 12: (Left:) Figure from Ijjas et al. [217] showing slow contraction in the multi-field case needed to generate scale invariant perturbations. Time evolution of the normalized energy densities in the fields (green) the curvature (red) and the shear (blue) are shown. The model drives the universe towards RW, but a non-zero gradient in the auxilliary field destabilizes the RW state on longer timescales. (Right:) Figure from Xue et al. [218] showing the local expansion as a function of space, evolving over time. Whilst the whole of the space bounces (going from negative to positive expansion), the space in the middle bounces at a much later time compared to the other regions, creating some inhomogeneity at late times.
A recent paper by Ijjas et al. [196] compared slow contraction and inflationary smoothing, concluding that the latter is more robust and thus strongly favoured. It is hard to assess such claims without running into philosophical questions about which models and which initial conditions are most likely to occur in nature (issues that include the measure problem mentioned above). The work also only considers the smoothing phase and does not take into account the difficulties of subsequently obtaining a bounce and generating primordial perturbations, which on the face of it would seem to favour inflation. Nevertheless, it provides a useful discussion of possible criteria for directly comparing two possible early universe models.
@@ 3.4 Signatures of inflationary bubble collisions
Figure 13: Schematic figure depicting the procedure from a scalar inflationary potential to a prediction of the observable comoving curvature perturbation $\mathcal{R}$. NR simulations are performed to construct the full collision spacetime (upper right). The perturbed RW metric inside the observation bubble is reconstructed by evolving a set of geodesics through the simulation (lower right panel). A gauge transformation is used to extract the comoving curvature perturbation $\mathcal{R}$ late in the inflationary epoch. Evolving the comoving curvature perturbation, the observable universe is split into a region that is, and a region that is not, affected by the collision. The lower left panel depicts the surface of last scattering inside the observation bubble in a reference frame where the observer is at the origin of coordinates. Figure from Wainwright et al. [242].
@@ 4.1 Reheating and oscillon formation
Another example of a preheating process is gauge preheating, when the energy of the inflaton is transferred into gauge fields. This can be sourced via a coupling $\alpha_{g}\phi F_{\mu\nu}\tilde{F}^{\mu\nu}$ in the Lagrangian, where $\alpha_{g}$ parameterizes the strength of the coupling and the range of momenta for which the process features a tachyonic instability. This instability allows the full vacuum energy to be transferred within a single oscillation of the inflaton field, and the strongest effects occur when the inflaton rolls the fastest ($\dot{\phi}$ is the largest) – typically at the end of inflation. In Adshead et al. [149], the authors showed that linearized gravity is quickly violated during this process. However, despite reaching large density contrasts $\delta\rho/\rho\approx 30$, PBHs did not form, indicating that the internal pressure of the gauge fields was playing an important role. Using a similar approach to Aurrekoetxea et al. [111] but based on the over-mass $\delta M$ enclosed in a region of radius $R$, they computed the compactness of the overdensities formed and found a maximum value of $\mathcal{C}\approx 0.1$. The authors compared the extracted GW spectrum from FLRW and BSSN lattice simulations finding remarkable agreement and suggesting that non-linear gravity at most had a $\mathcal{O}(1)$ effect, see Fig. 17. Finally, we note that although motivated by dark matter, the work of Alcubierre et al. [258] showing the growth of scalar fluctuations on an expanding background is also relevant to this section, as is the early work on nucleosynthesis in 1D planar symmetry by Centrella et al. [17], which showed that inhomogeneities affect predicted nucleosynthesis abundances.
@@ 4.2 Phase transitions and topological defects
As the universe expands and cools down, it can undergo phase transitions that lead to the formation of exotic objects. Among the most studied relics are topological defects [259, 260, 261] – structures formed when the universe experienced symmetry breaking. The specific details of these structures depend on the topology of the vacuum. For example, if the vacuum breaks a parity symmetry, two-dimensional defects known as domain walls can form. We previously encountered such relics in Sec. 3.4, where we discussed the observational signatures in the CMB from inflationary bubble collisions. NR has also been used to study the evolution of individual bubbles during vacuum decay, such as the 1+1D work by Blanco-Pillado et al. [262], and Giombi et al. [263]. In the context of the Standard Model, East et al. [86] studied the evolution of unstable fluctuations in the Higgs field during inflation using numerical relativity simulations in axisymmetry. Recently, Lin et al. [264, 265] have performed 3+1D NR simulations to investigate how the evolution of inhomogeneities in the early universe can populate the landscape of multiple vacua. The dynamics of the formed bubbles may serve as seeds for supermassive black holes, as studied by Huang et al. [266], and enhance the formation of black hole binaries, as suggested by Yuwen et al. [267].
@@
The dynamics of topological defects sources gravitational waves. If these are weak, we might observe the incoherent sum of all the events that are too weak to be directly detected – the stochastic GW background. However, local dynamics might also source strong gravitational-wave signatures that can be directly detected. In both cases, an NR treatment may be necessary: either to validate the predictions of a weak-field treatment of the source, or to predict the detailed signatures to be expected. However, much of the current research focuses on the linearised treatment of these gravitational-wave signatures, and to the best of our knowledge, there is no simulation that fully incorporates NR in a cosmological context – hence the brevity of this section. However, asymptotically flat NR simulations have been performed to understand the GR dynamics of these objects and construct gravitational-wave templates [270, 271, 272, 273, 274].
The dynamics of bubbles after a cosmological phase transition has also recently been studied using holography. In this approach, one solves Einstein’s equations in asymptotically AdS5 to study the out-of-equilibrium dynamics of strongly-coupled gauge theories on the boundary [275, 276, 277, 278].
@@ 4.3 Primordial black holes
Using spheroidal super-horizon perturbations, Yoo et al. [320] studied the effect of ellipticity on the threshold of PBH formation, finding that it is negligible in standard situations during the radiation-dominated era. Yoo later extended the work [321] by adding a tidal torque to the non-spherical profile of the initial curvature perturbation as a mechanism to generate angular momentum and form a spinning remnant, but the final spin remained consistent with zero, in agreement with expectations for the radiation-dominated era. More recently, Escrivá & Yoo [322] have used 3+1 relativistic simulations to study the collapse of ellipsoidal perturbations and found that while the BH formation threshold is modified, the modification is small. However, in the case of a softer equation of state, small deviations from spherical symmetry may evolve differently due to the absence of pressure gradient forces. De Jong et al. [323, 112] used a simple model of two scalar fields to study the mass and spin of PBHs from the collapse of perturbations with angular momentum during an early matter-dominated era $P\approx 0$. They found that although the initial mass and spin at formation can be of order $\mathcal{O}(10\%)$, the accretion of background matter post-collapse dilutes the dimensionless spin of the PBH. Hence, unless the matter era is short, they concluded that the final spin of the PBHs formed will be negligible.